Bramble Company uses a periodic inventory system. For April, when the company sold 650 units, the following information is available.

Units

Unit Cost

Total Cost

April 1 inventory

350 $24 $ 8,400

April 15 purchase

370 29 10,730

April 23 purchase

280

31

8,680

1,000

$27,810


Compute the April 30 inventory and the April cost of goods sold using the LIFO method.
Answer #1

Inventory = 350*24

= 8,400

Cost of Goods Sold using Lifo

= (280*31) + (370*29)

= 8680 + 10730

= 19,410